OpenStudy (anonymous):

Slope of the curve y^3-xy^2=4 at point where y=2 is ____

OpenStudy (amistre64):

what our implicit derivative?

OpenStudy (anonymous):

y^3-y^2=0?

OpenStudy (amistre64):

hmm, not quite y^3 power rule it: 3y^2 y' -xy^2 is a product so .. -x'y^2 -2xy y' and 4 is a constant so it goes to 0

OpenStudy (amistre64):

solve for y', and you need to solve for x as well, then all the parts are there to plug in

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

3y^2+2x'y^3y'=0

OpenStudy (anonymous):

y'=-3/2x'y

OpenStudy (amistre64):

y^3 - xy^2 = 4 3y^2 y' -x'y^2 - 2xy y' = 0 let y=2 2^3 - x 2^2 = 4 3 2^2 y' -2^2 - 2x 2 y' = 0 ------ 8 - 4x = 4 , x=1 12y' -4 - 4y' = 0 , now all we need is to solve for y'

OpenStudy (amistre64):

most times its simpler to just solve for y' after all the other parts are in place

OpenStudy (anonymous):

i think i understand

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

but wait

OpenStudy (anonymous):

what happened to the x'

OpenStudy (amistre64):

x', with respect to x is just: dx/dx which is equal to 1

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

what is the rate of change of x as x moves by 1? well, 1 of course :)

OpenStudy (amistre64):

now if we had x as a function of time, then dx/dt would have to have been given to us or left alone as x'

OpenStudy (anonymous):

i think that is one of the later problems on my sheet

OpenStudy (anonymous):

y'=-1/2

OpenStudy (anonymous):

thank you a TON!

OpenStudy (amistre64):

12y' -4 - 4y' = 0 12y' - 4y' = 4 8y' = 4 , im getting +1/2

OpenStudy (anonymous):

oh whoops forgot the negative

OpenStudy (anonymous):

wow that was bad

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

:)